# Photon Statistics for Fock and Coherent States Interfering in a Beamsplitter

### v1

@@ 3.5 Special Cases
where $\mathcal{D}(\xi)=\exp(\xi a^{\dagger}-\xi^{*}a)$ is the displacement operator for complex amplitude $\xi$.
The BS transforms coherent states into coherent states. Using the transformation property:
@@
This shows the BS linearly transforms the input amplitudes $(\alpha,\beta)$ to output amplitudes $(t\alpha+ir\beta,ir\alpha+t\beta)$.
@@ 4.3 Mandel Q parameter
$Q_{c}=Q_{d}=0,$ (4.7)
confirming Poissonian statistics are preserved.
@@ 4.5 Special Cases
We recover equation (3.12) exchanging $t$ for $r.$
@@ 5 Conclusions
In this work, we have presented a theoretical analysis of the photon statistics resulting from the interference of quantum light fields at an ideal lossless beamsplitter (BS). By exploring a broad range of input states, namely Fock–Fock, Fock–coherent, and coherent–coherent configurations, we derived explicit analytical expressions for the output states and computed fundamental statistical quantities, including the mean photon number, photon number variance, Mandel $Q$ parameter, and $g$ function. These calculations were carried out using standard operator techniques and the well-known properties of coherent and Fock states, which highlight the quantum structure underlying the field transformations induced by the BS.
@@
Beyond these specific cases, our study provides a unified perspective on how different statistical regimes, ranging from purely quantum (sub-Poissonian) to classical-like (Poissonian), emerge from the linear mixing of quantum states at a BS. This viewpoint is particularly relevant for understanding fundamental experiments in quantum optics, such as single-photon interference and coherent state splitting, as well as for interpreting modern applications in quantum information processing and photonic technologies.
@@ Acknowledgements.
J. A. T. S. is grateful for the encouragement of Prof. Danieverton Moretti and the financial support provided by UFCG.

### v2

@@ 3.5 Special Cases
where $\mathcal{D}(\xi)=\exp(\xi a^{\dagger}-\xi^{*}a)$ is the displacement operator for complex amplitude $\xi$, and $|\alpha\rangle$ and $|\beta\rangle$ are coherent states with amplitudes $\alpha$ and $\beta$, respectively. Coherent states are quantum states that closely resemble classical harmonic oscillators, exhibiting minimum uncertainty and Poissonian photon statistics [1]. Using the transformation property:
@@
This means that coherent states remain coherent after passing through the BS, with their amplitudes linearly mixed according to the BS coefficients $t$ and $r$. Importantly, the linearity of the BS ensures that no entanglement is generated between the output modes when both inputs are classical coherent states. This illustrates the close correspondence between quantum coherent states and classical electromagnetic waves, particularly in linear optical systems.
@@ 4.1 Mean Photon Number
This behavior aligns with classical wave theory, where the electric field amplitudes of light waves combine linearly, and the resulting intensity fluctuations follow a Poisson distribution [27, 28].
@@ 4.3 Mandel Q parameter
$Q_{c}=Q_{d}=0.$ (4.7)
This result emphasizes the classical nature of coherent states under linear optics. Since both outputs remain coherent and separable, no entanglement or nonclassical correlations are generated. This property underlies the correspondence between classical interference patterns in optics and their quantum descriptions when restricted to coherent states.
@@ 4.5 Special Cases
We recover equation (3.13) exchanging $t$ for $r$, given that $|\alpha=0\rangle$ is the same as that of $|n=0\rangle$.
@@ 5 Conclusions
In this work, we have presented a theoretical analysis of the photon statistics resulting from the interference of quantum light fields at an ideal, lossless beamsplitter (BS). By exploring a broad range of input states, namely Fock–Fock, Fock–coherent, and coherent–coherent configurations, we derived explicit analytical expressions for the output states and computed fundamental statistical quantities, including the mean photon number, photon number variance, Mandel $Q$ parameter, and $g$ function. These calculations were carried out using standard operator techniques and the well-known properties of coherent and Fock states, which highlight the quantum structure underlying the field transformations induced by the BS.
@@
Beyond these specific cases, our study provides a unified perspective on how different statistical regimes, ranging from purely quantum (sub-Poissonian) to classical-like (Poissonian), emerge from the linear mixing of quantum states at a BS. This viewpoint is particularly relevant for understanding fundamental experiments in quantum optics, such as single-photon interference and coherent state splitting, as well as for interpreting modern applications in quantum information processing and photonic technologies. However, we note that the model considered here is intentionally simplified and didactic. In practical implementations, additional degrees of freedom, especially photon polarization and the spectral (frequency) dependence of the field operators, $a=a(\omega)$, play a crucial role in determining the observed interference and photon statistics. For a detailed treatment of these effects in multi-photon interference and their application to measurement-device-independent quantum key distribution, see [43].
@@ Acknowledgements.
The author gratefully acknowledges the encouragement of Prof. Danieverton Moretti, the constructive criticisms and suggestions of the anonymous referees, and the financial support provided by UFCG.